Suppose x is distributed with a mean of 25.1 and a standard deviation of 3.2. What is the probability of randomly selecting an individual from the distribution with value that is greater than 21.2 and less than 23.4
1
Expert's answer
2020-10-27T19:00:13-0400

"\mu=25.1"

"\sigma=3.2"

"P(21.2<X<23.4)=P(\frac{21.2-\mu}{\sigma}<Z<\frac{23.4-\mu}{\sigma})="

"=P(\frac{21.2-25.1}{3.2}<Z<\frac{23.4-25.1}{3.2})=P(-1.22<Z<-0.53)="

"=P(Z<-0.53)-P(Z<-1.22)=0.2981-0.1112=0.1869"


Answer: 18.69% of individuals are selected between 21.2 and 23.4.